a women invest $25,000 in two business ventures/ last year she made a profit of 15% from the first venture but lost 5% from the second venture. if last year's income from the two investments was equivalent to a return of 8% of the entire amount invested, how much had she invested in each venture?
Answer by mananth(16946)

a women invest $25,000 in two business ventures/ last year she made a profit of 15% from the first venture but lost 5% from the second venture. if last year's income from the two investments was equivalent to a return of 8% of the entire amount invested, how much had she invested in each venture?
Let x be invested at 15%
25000-x is invested at 5%
Income = 8% of 25000 -------0.08*25000
0.15x-0.05(25000-x)=0.08*25000
0.15x -1250+0.05x=2000
0.20 x= 2000+1250
0.20x= 3250
x=16,250 @ 15%
25000-7500 = 8750 @ 5%
$16,250 at 15%
$ 8,750 at 5%